# The discrete generalized exchange-driven system

### v1

@@ 5.1. The isolated case (7)
In this section we formulate and prove the existence of solutions for the initial value problem (7), (8) in the isolated system case.
Theorem 5.1.
Let $c_{0}=(c_{0i})_{i\geq 0}\in X_{0,1}^{+}$. Assume the rate coefficients satisfy the conditions (27)–(28). Then the initial value problem (7), (8) for the isolated case has at least one global solution $c$ defined on the interval $[0,\infty)$.
Proof. By using a refined version of de la Vallée-Poussin theorem as in [18, Section 4], we can conclude that our hypothesis for $c_{0}$ implies that there is $\sigma_{0}\in\mathcal{E}$ such that
@@
where :=$\mathcal{I}:=\{(i,j,k)\in({\mathbb{N}}_{+})^{3}:k\leqslant i\},$ with domain $D_{\Phi}$ made by the sequences $\beta\in X_{0,1}^{+}$ for which the above triple series is convergent. We prove that, in fact, $D_{\Phi}=X_{0,1}^{+}.$ Let $\eta\in{\mathbb{N}}_{+}.$ Then, for each $\beta\in X_{0,1}^{+},$ by (27) and (28), we have,
@@
[…]
@@
[…]
$\mathcal{J}^{N}_{2,i}$ :=$:=\{(p,q)\in{\mathbb{N}}^{2}:1\leqslant p\leqslant N,\,q=i\},\quad i\in{\mathbb{N}}_{+},$
$\mathcal{J}^{N}_{3,i}$ :=$:=\{(p,q)\in{\mathbb{N}}^{2}:1\leqslant q\leqslant i-1,\,i-q\leqslant p\leqslant N\},\quad i\in{\mathbb{N}}_{2},$
[…]
@@
$\mathcal{K}^{N}_{1,i}(p,q)$ :=$:=\{p-i\},\quad i\in{\mathbb{N}},$
$\mathcal{K}^{N}_{2,i}(p,q)$ :=$:=\{k\in{\mathbb{N}}:1\leqslant k\leqslant\min(p,N-i)\},\quad i\in{\mathbb{N}}_{+},$
$\mathcal{K}^{N}_{3,i}(p,q)$ :=$:=\{i-q\},\quad i\in{\mathbb{N}}_{2},$
$\mathcal{K}^{N}_{4,i}(p,q)$ :=$:=\{k\in{\mathbb{N}}:1\leqslant k\leqslant\min(N-q,i)\},\quad i\in{\mathbb{N}}_{+}.$
@@
$\begin{split}|\hat{Q}_{j,i}^{N}(c^{N})-\hat{Q}_{j,i}(c)|_{L^{1}(0,T)}&\\ &\leqslant\sum_{(p,q)\in\mathcal{J}^{M}_{j,i}}\;\sum_{k\in\mathcal{K}^{M}_{j,i}(p,q)}a(p,q;k)\big{|}c^{N}_{p}c^{N}_{q}-c_{p}c_{q}\big{|}_{L_{1}(0,T)}\\ &\qquad+\Big{|}\sum_{(p,q)\in\mathcal{J}^{N}_{j,i}\setminus\mathcal{J}^{M}_{j,i}}\;\sum_{k\in\mathcal{K}^{N}_{j,i}(p,q)}a(p,q;k)c^{N}_{p}c^{N}_{q}\Big{|}_{L_{1}(0,T)}\\ &\qquad+\Big{|}\sum_{(p,q)\in\mathcal{J}^{N}_{j,i}\setminus\mathcal{J}^{M}_{j,i}}\;\sum_{k\in\mathcal{K}^{N}_{j,i}(p,q)}a(p,q;k)c_{p}c_{q}\Big{|}_{L_{1}(0,T)}\end{split}$ (89)
@@
$\mathcal{J}^{N}_{4,i}\setminus\mathcal{J}^{M}_{4,i}$ =$=\{i\}\times\big{(}[M,N-1]\cap{\mathbb{N}}\big{)},$
@@
[…]
$J_{II}$ :=$:=\{(p,q)\in{\mathbb{N}}^{2}:i+1\leqslant p\leqslant M,\;M+1+i-p\leqslant q\leqslant N+i-p\}.$
@@
$\mathcal{I}_{1,i}^{N}$ :=$:=\{(p,k)\in{\mathbb{N}}^{2}:i+1\leqslant p\leqslant N,\;k=p-i\},\quad\,i\in\{1,\dots,N-1\},$
[…]
$\mathcal{I}_{3,i}^{N}$ :=$:=\{(p,k)\in{\mathbb{N}}^{2}:i+1\leqslant p\leqslant N,\;k=i\},\quad\qquad i\in\{1,\dots,N\}.$
@@
[…]
@@
$\mathcal{I}_{1,i}^{N}\setminus\mathcal{I}_{1,i}^{M}$ =$=\{(p,k)\in{\mathbb{N}}^{2}:M+1\leqslant p\leqslant N,\;k=p-i\},\quad i\in\{1,\dots,N-1\},$
[…]
$\mathcal{I}_{3,i}^{N}\setminus\mathcal{I}_{3,i}^{M}$ =$=\{(p,k)\in{\mathbb{N}}^{2}:M+1\leqslant p\leqslant N,\;k=i\},\quad i\in\{1,\dots,N\}.$
@@
Theorem 5.2.
and
Proof.
Let $T>0$ and $2\leqslant L<N$. Using Corollary 4.2 we write, for all $t\in[0,T).$
@@
and now repeating the process of letting first $N\to\infty$, then $L\to\infty$, and using (85) and (31), we conclude (112).∎
@@ 5.2. The non-isolated case
In the non-isolated case theorems 5.1 and 5.2 can be stated almost exactly verbatim (the only difference being the lack of conservation of the cluster number density) with essentially the same proofs, where the only difference is that the terms $Q_{j,0}$ are now all zero. For completeness we state them now, omitting the proofs.
Theorem 5.3.
Let $c_{0}=(c_{0i})_{i\geq 0}\in X_{0,1}^{+}$. Assume the rate coefficients satisfy the conditions (27)–(28). Then the initial value problem (7), (8) for the non-isolated case has at least one global solution $c$ defined on the interval $[0,\infty)$.
Theorem 5.4.
@@ 6. A regularity result
Theorem 6.1.
Suppose that the rate coefficients satisfy (27)-(28), for all $j\in{\mathbb{N}}.$ If $c=(c_{i})$ is a mild solution of (7)-(8) on $[0,T)$, in either the isolated or the non-isolated cases, and if $\sum_{i=1}^{\infty}ic_{i}(\cdot)$ is constant in $[0,T),$ then, for each $i\in{\mathbb{N}},$ $c_{i}\in C^{1}([0,T),\mathbb{R}^{+}_{0}).$
Proof.
Similar to the function $\Phi$ introduced in the proof of Theorem 5.1, we define the function $\Psi$ by
@@
This shows that, in fact, $D_{\Psi}=X^{+}_{0,1}.$ On the other hand, from the assumption that $\sum_{i=1}^{\infty}ic_{i}(t)$ is constant for $t\in[0,T)$ Dini’s theorem allow us to conclude that this series of nonnegative functions is uniformly convergent on any compact subset of $[0,T)$. And so $\sum_{i=0}^{\infty}c_{i}(t)$ is also uniformly convergent on compact subsets of $[0,T)$. This, together with (114) ensure that the same is true for the series defining $\Psi(c(\cdot))$. Since, for $j=1,2,3,4$ with $i\in\mathbb{N}_{+}$, and $j=1,2$ with $i=0$, the the indices of sums defining $Q_{j,i}(c)$ run over subsets of $\mathcal{I}_{0},$ we conclude that these are also uniformly convergent on any compact subset of $[0,T)$ and therefore, by the continuity of $c,$ we conclude that $Q_{j,i}(c(\cdot))$ is continuous in $[0,T)$. But then, by (iii) in Definition 3.1, the thesis of the lemma is obtained. ∎
@@ 7. Uniqueness of solutions
In this section, we establish and prove a partial uniqueness result for (7)–(8) under some restrictions on the reaction rate. The proof is based on ideas commonly used in studies of coagulation-type equations (see, e.g., [1, 18]) and so we will skip most of the details of the computations. Those ideas are the following: we assume that, for some $T\in(0,\infty)$ the initial value problem (7)–(8) has two solutions $c=(c_{i})$ and $d=(d_{i})$. Defining $x:=c-d$, we prove that, on $[0,T]$, this function satisfies a differential inequality
[…]
@@
Let $c=(c_{i})_{i\geqslant 0}$ be any solution of (7) in the isolated case. Then, for every $n\in\ N$ and sequence $(g_{i})$ we have
@@
Observe that in (115), when $k=n$ the only term of the triple sum has $(i,j,k)=(n,0,n)$ and so, either from $g_{j+k}+g_{i-k}-g_{j}-g_{i}=g_{n}+g_{0}-g_{0}-g_{n}=0$, or by (9), its contribution to (115) is identically zero. This means that the sum in $k$ can be done just from $1$ to $n-1$, exactly as in (43).
@@
We can now state the following uniqueness result:
Theorem 7.2.
Proof.
Suppose $c=(c_{i})$ and $d=(d_{i})$ are two solutions of (7)–(8) on $[0,T].$ Let $x:=c-d$. Note that, since that $c(0)$ and $d(0)$ are both equal to the initial condition $c_{0}$, and thus $x(0)=0$. Also, for all $i$ and $j$ it is clear that $c_{i}c_{j}-d_{i}d_{j}=c_{i}x_{j}+d_{j}x_{i}.$ Observing that, for every absolutely continuous function $u(t)$, the function $t\mapsto|u(t)|$ is also absolutely continuous and $\frac{d}{dt}|u(t)|=\operatorname{sgn}(u(t))\frac{du}{dt}$ almost everywhere (where $\operatorname{sgn}(\cdot)$ denotes the sign function), we can use Proposition 7.1 to write
@@
where $\tilde{g}(i,j;k):=g_{j+k}+g_{i-k}-g_{j}-g_{i}$, with $g_{i}=(1+i^{\alpha})\operatorname{sgn}(x_{i})$, and $\Delta R_{j,n}(\cdot):=R_{j,n}(c(\cdot))-R_{j,n}(d(\cdot)).$ Writing $x_{\ell}=\operatorname{sgn}(x_{\ell})|x_{\ell}|$ and remembering that the sign function is either $-1,0$ or $1$, we can write
@@
whence $x_{i}(t)\equiv 0$, and the uniqueness is proved. ∎

### v2

@@ 5.1. The isolated case (7)
In this section prove the existence of solutions for the initial value problem (7), (8) in the isolated system case.
Proof of theorem 3.2: By using a refined version of de la Vallée-Poussin theorem as in [18, Section 4], we can conclude that our hypothesis for $c_{0}$ implies that there is $\sigma_{0}\in\mathcal{E}$ such that
@@
where $\mathcal{I}:=\{(i,j,k)\in({\mathbb{N}}_{+})^{3}:k\leqslant i\},$ with domain $D_{\Phi}$ made by the sequences $\beta\in X_{0,1}^{+}$ for which the above triple series is convergent. We prove that, in fact, $D_{\Phi}=X_{0,1}^{+}.$ Let $\eta\in{\mathbb{N}}_{+}.$ Then, for each $\beta\in X_{0,1}^{+},$ by (27) and (28), we have,
@@
$\leqslant\;$ $C\mathcal{Q}\Big{(}\sum_{i=1}^{\eta}i\beta_{i}\Big{)}\Big{(}\sum_{j=1}^{\eta}j\beta_{j}\Big{)}+C\mathcal{Q}\Big{(}\sum_{i=1}^{\eta}i\beta_{i}\Big{)}\Big{(}\sum_{j=1}^{\eta}\beta_{j}\Big{)}$
$\leqslant\;$ $2C\mathcal{Q}\|\beta\|^{2}.$
@@
$\mathcal{J}^{N}_{1,i}$ $:=\{(p,q)\in{\mathbb{N}}^{2}:i+1\leqslant p\leqslant N,\,1\leqslant q\leqslant N+i-p\},\quad i\in{\mathbb{N}},$
$\mathcal{J}^{N}_{2,i}$ $:=\{(p,q)\in{\mathbb{N}}^{2}:1\leqslant p\leqslant N,\,q=i\},\quad i\in{\mathbb{N}}_{+},$
$\mathcal{J}^{N}_{3,i}$ $:=\{(p,q)\in{\mathbb{N}}^{2}:1\leqslant q\leqslant i-1,\,i-q\leqslant p\leqslant N\},\quad i\in{\mathbb{N}}_{2},$
$\mathcal{J}^{N}_{4,i}$ $:=\{(p,q)\in{\mathbb{N}}^{2}:p=i,\,1\leqslant q\leqslant N-1\},\quad i\in{\mathbb{N}}_{+},$
@@
$\mathcal{K}^{N}_{1,i}(p,q)$ $:=\{p-i\},\quad i\in{\mathbb{N}},$
$\mathcal{K}^{N}_{2,i}(p,q)$ $:=\{k\in{\mathbb{N}}:1\leqslant k\leqslant\min(p,N-i)\},\quad i\in{\mathbb{N}}_{+},$
$\mathcal{K}^{N}_{3,i}(p,q)$ $:=\{i-q\},\quad i\in{\mathbb{N}}_{2},$
$\mathcal{K}^{N}_{4,i}(p,q)$ $:=\{k\in{\mathbb{N}}:1\leqslant k\leqslant\min(N-q,i)\},\quad i\in{\mathbb{N}}_{+}.$
@@
$\begin{split}|\hat{Q}_{j,i}^{N}(c^{N})-\hat{Q}_{j,i}(c)|_{L^{1}(0,T)}&\\ &\hskip-28.45274pt\leqslant\sum_{(p,q)\in\mathcal{J}^{M}_{j,i}}\;\sum_{k\in\mathcal{K}^{M}_{j,i}(p,q)}a(p,q;k)\big{|}c^{N}_{p}c^{N}_{q}-c_{p}c_{q}\big{|}_{L_{1}(0,T)}\\ &\hskip-28.45274pt\qquad+\Big{|}\sum_{(p,q)\in\mathcal{J}^{N}_{j,i}\setminus\mathcal{J}^{M}_{j,i}}\;\sum_{k\in\mathcal{K}^{N}_{j,i}(p,q)}a(p,q;k)c^{N}_{p}c^{N}_{q}\Big{|}_{L_{1}(0,T)}\\ &\hskip-28.45274pt\qquad+\Big{|}\sum_{(p,q)\in\mathcal{J}^{N}_{j,i}\setminus\mathcal{J}^{M}_{j,i}}\;\sum_{k\in\mathcal{K}^{N}_{j,i}(p,q)}a(p,q;k)c_{p}c_{q}\Big{|}_{L_{1}(0,T)}\end{split}$ (91)
@@
$\mathcal{J}^{N}_{4,i}\setminus\mathcal{J}^{M}_{4,i}$ $=\{i\}\times\big{(}[M,N-1]\cap{\mathbb{N}}\big{)},$
@@
$J_{I}$ $:=\{(p,q)\in{\mathbb{N}}^{2}:M+1\leqslant p\leqslant N,\;1\leqslant q\leqslant N+i-p\},$
$J_{II}$ $:=\{(p,q)\in{\mathbb{N}}^{2}:i+1\leqslant p\leqslant M,\;M+1+i-p\leqslant q\leqslant N+i-p\}.$
@@
$\mathcal{I}_{1,i}^{N}$ $:=\{(p,k)\in{\mathbb{N}}^{2}:i+1\leqslant p\leqslant N,\;k=p-i\},\quad\,i\in\{1,\dots,N-1\},$
$\mathcal{I}_{2,0}^{N}$ $:=\{(p,k)\in{\mathbb{N}}^{2}:2\leqslant p\leqslant N,\;1\leqslant k\leqslant p-1\},$
$\mathcal{I}_{3,i}^{N}$ $:=\{(p,k)\in{\mathbb{N}}^{2}:i+1\leqslant p\leqslant N,\;k=i\},\quad\qquad i\in\{1,\dots,N\}.$
@@
$\mathcal{J}^{N}:=\{(p,k)\in({\mathbb{N}}_{+})^{2}:1\leqslant p\leqslant N,\;1\leqslant k\leqslant p-1\},$
@@
$\mathcal{I}_{1,i}^{N}\setminus\mathcal{I}_{1,i}^{M}$ $=\{(p,k)\in{\mathbb{N}}^{2}:M+1\leqslant p\leqslant N,\;k=p-i\},\quad i\in\{1,\dots,N-1\},$
$\mathcal{I}_{2,0}^{N}\setminus\mathcal{I}_{2,0}^{M}$ $=\{(p,k)\in{\mathbb{N}}^{2}:M+1\leqslant p\leqslant N,\;1\leqslant k\leqslant p-1\},$
$\mathcal{I}_{3,i}^{N}\setminus\mathcal{I}_{3,i}^{M}$ $=\{(p,k)\in{\mathbb{N}}^{2}:M+1\leqslant p\leqslant N,\;k=i\},\quad i\in\{1,\dots,N\}.$
@@
Proof of theorem 3.3: Let $T>0$ and $2\leqslant L<N$. Using Corollary 4.2 we write, for all $t\in[0,T).$
@@
and now repeating the process of letting first $N\to\infty$, then $L\to\infty$, and using (87) and (33), we conclude (32). $\square$
@@ 5.2. The non-isolated case
In the non-isolated case theorems 3.2 and 3.3 can be stated almost exactly verbatim (the only difference being the lack of conservation of the cluster number density) with essentially the same proofs, where the only difference is that the terms $Q_{j,0}$ are now all zero. For completeness we state them now, omitting the proofs.
@@ 6. A regularity result
Proof of the theorem 3.6: Similar to the function $\Phi$ introduced in the proof of Theorem 3.2, we define the function $\Psi$ by
@@
This shows that, in fact, $D_{\Psi}=X^{+}_{0,1}$. On the other hand, from the assumption that $\sum_{i=1}^{\infty}ic_{i}(t)$ is constant for $t\in[0,T)$ Dini’s theorem allow us to conclude that this series of nonnegative functions is uniformly convergent on any compact subset of $[0,T)$. And so $\sum_{i=0}^{\infty}c_{i}(t)$ is also uniformly convergent on compact subsets of $[0,T)$. This, together with (114) ensure that the same is true for the series defining $\Psi(c(\cdot))$. Since, for $j=1,2,3,4$ with $i\in\mathbb{N}_{+}$, and $j=1,2$ with $i=0$, the the indices of sums defining $Q_{j,i}(c)$ run over subsets of $\mathcal{I}_{0},$ we conclude that these are also uniformly convergent on any compact subset of $[0,T)$ and therefore, by the continuity of $c,$ we conclude that $Q_{j,i}(c(\cdot))$ is continuous in $[0,T)$. But then, by (iii) in Definition 3.1, the thesis of the lemma is obtained. $\square$
@@ 7. Uniqueness of solutions
In this section, we prove the partial uniqueness result for (7)–(8) stated in theorem 3.7. The proof is based on ideas commonly used in studies of coagulation-type equations (see, e.g., [1, 18]) and so we will skip most of the details of the computations. Those ideas are the following: we assume that, for some $T\in(0,\infty)$ the initial value problem (7)–(8) has two solutions $c=(c_{i})$ and $d=(d_{i})$. Defining x:=cdx:=c-d, we prove that, on $[0,T]$, this function satisfies a differential inequality
$\psi(t)\leqslant\int_{0}^{t}\psi(s)ds,$
@@
Let $c=(c_{i})_{i\geqslant 0}$ be any solution of (7) in the isolated case. Then, for every $n\in{\mathbb{N}}$ and sequence $(g_{i})$ we have
@@
Observe that in (115), when $k=n$ the only term of the triple sum has $(i,j,k)=(n,0,n)$ and so, either from $g_{j+k}+g_{i-k}-g_{j}-g_{i}=g_{n}+g_{0}-g_{0}-g_{n}=0$, or by (9), its contribution to (115) is identically zero. This means that the sum in $k$ can be done just from $1$ to n1n-1, exactly as in (45).
@@
We can now prove the uniqueness result stated in theorem 3.7.
Proof of the theorem 3.7: Suppose $c=(c_{i})$ and $d=(d_{i})$ are two solutions of (7)–(8) on $[0,T].$ Let x:=cdx:=c-d. Note that, since that $c(0)$ and $d(0)$ are both equal to the initial condition $c_{0}$, and thus $x(0)=0$. Also, for all $i$ and $j$ it is clear that $c_{i}c_{j}-d_{i}d_{j}=c_{i}x_{j}+d_{j}x_{i}.$ Observing that, for every absolutely continuous function $u(t)$, the function $t\mapsto|u(t)|$ is also absolutely continuous and $\frac{d}{dt}|u(t)|=\operatorname{sgn}(u(t))\frac{du}{dt}$ almost everywhere (where $\operatorname{sgn}(\cdot)$ denotes the sign function), we can use Proposition 7.1 to write
@@
where $\tilde{g}(i,j;k):=g_{j+k}+g_{i-k}-g_{j}-g_{i}$, with $g_{i}=(1+i^{\alpha})\operatorname{sgn}(x_{i})$, and $\Delta R_{j,n}(\cdot):=R_{j,n}(c(\cdot))-R_{j,n}(d(\cdot)).$ Writing $x_{\ell}=\operatorname{sgn}(x_{\ell})|x_{\ell}|$ and remembering that the sign function is either 1,0-1,0 or $1$, we can write
@@
whence $x_{i}(t)\equiv 0$, and the uniqueness is proved. $\square$